The hypotenuse of a right triangle is 25 inches. Find the length of the legs if their sum is 31 inches.
Answer by edjones(8007)

c=25 a+b=31
b=31-a
a^2+b^2=c^2
a^2+(31-a)^2=625
a^2+a^2-62a+961=625
2a^2-62a+336=0
a^2-31a+168=0
(a-24)(a-7)=0
a=24, b=7 or vice versa
.
Ed